### v1

# MAC-VO: Metrics-aware Covariance for Learning-based Stereo Visual Odometry mac-vo.github.io

Abstract
We propose the MAC-VO, a novel learning-based stereo VO that leverages the learned metrics-aware matching uncertainty for dual purposes: selecting keypoint and weighing the residual in pose graph optimization. Compared to traditional geometric methods prioritizing texture-affluent features like edges, our keypoint selector employs the learned uncertainty to filter out the low-quality features based on global inconsistency. In contrast to the learning-based algorithms that model the scale-agnostic diagonal weight matrix for covariance, we design a metrics-aware covariance model to capture the spatial error during keypoint registration and the correlations between different axes. Integrating this covariance model into pose graph optimization enhances the robustness and reliability of pose estimation, particularly in challenging environments with varying illumination, feature density, and motion patterns. On public benchmark datasets, MAC-VO outperforms existing VO algorithms and even some SLAM algorithms in challenging environments. The covariance map also provides valuable information about the reliability of the estimated poses, which can benefit decision-making for autonomous systems.
Index Terms:
SLAM, Learning VO, Covariance Estimation
I Introduction
Visual Odometry (VO) predicts the relative camera pose from image sequences and often serves as the front-end of Simultaneous Localization and Mapping (SLAM) systems. Over the past few decades, both geometric and learning-based methods have been developed with significant advances in generalizability and accuracy [1, 2, 3, 4]. However, VO remains a challenging problem in real-world scenarios, with multiple visual degraded scenarios such as low illumination, dynamic and texture-less scenes.
To improve the robustness in challenging scenes, geometric-based VO algorithms employ outlier filtering strategies [5] and weigh the optimization residuals by the covariance of the observed features [6]. However, how to effectively select the reliable keypoints and model their covariance becomes two significant challenges. Existing methods typically select the keypoints based on local intensity gradient with a manually defined threshold [7, 8, 9]. These approaches leads to errors and outliers because it doesn’t model the structure or context information of the environment (e.g. features on repetitive patterns may not be ideal candidates despite high image gradients). Moreover, the covariance model is often empirically modeled using a constant parameter, which is sub-optimal. In addition, the parameters in keypoint selection and covariance model need to be extensively tuned for different environments.
Figure 1: Without relying on multi-frame bundle adjustment, MAC-VO aggregately reconstruct the map based on the two-frame Pose graph optimization. We propose metrics-aware covariance model for 3D keypoint based on the learning matching uncertainty. The covariance model serves dual purposes: selecting keypoints and weighing the residual in the pose graph optimization.
With advances in learning-based visual features, more algorithms utilize learned features [7] to optimize the camera pose. Confidence score [10] or confidence weights [3, 4] of these feature points are often obtained in an unsupervised manner. The learned confidence helps to track the features and model the reliability during the optimization. However, these confidence or uncertainty values are scale-agnostic, which means they don’t reflect the actual estimation error in the 3D space. This scale-agnostic problem brings two limitations. Firstly, it makes the covariance inconsistent across different environments that vary in scale, such as indoors and outdoors. Secondly, it makes it harder to integrate multiple constraints from different modalities or sensors.
To overcome the above challenges, this paper addresses the problem of modeling metrics-aware covariance values for the 3D keypoints. More specifically, this is achieved through two innovations. Firstly, we propose a learning-based model to quantify the 2D metrics-aware uncertainty of feature matching. Inspired by the FlowFormer [11] and GMA [12], we employ an iterative update model and motion aggregator to predict the uncertainty in 2D image space, which helps to filter unreliable features in the occluded region or low-illumination area. Secondly, based on the learned 2D uncertainty values, we model the covariance of the feature points in 3D space using a metrics-aware 3D covariance model. Compared to DROID-SLAM [3], which utilizes a scale-agnostic diagonal covariance matrix, our approach provides a more accurate representation by modeling the covariance of 3D feature points. This covariance model includes the inter-axes correlation of the 3D features. In the ablation study, we demonstrate the inter-frame consistency and the intra-frame consistency of our proposed covariance model.
We integrate the above two innovations into MAC-VO, a stereo VO that features superior keypoint selection and pose graph optimization based on the metrics-aware covariance model and achieves accurate tracking results in challenging cases compared with state-of-the-art VOs and even some SLAM systems without fine-tuning and without multi-frame optimization. In summary, the main contributions are:
  • We present a learning-based 2D uncertainty network with metrics awareness, leveraging iterative motion aggregator to capture the inconsistency of the feature matching. This metrics-aware uncertainty evaluates the quality of features in the keypoint selector and weights the residual of backend optimization.
  • This paper introduces a novel metrics-aware 3D covariance model based on the 2D uncertainty of feature matching and depth estimation. The ablation study demonstrates the necessity of scale consistency and the off-diagonal terms in the pose graph optimization.
  • We propose the MAC-VO, a stereo VO pipeline that estimates the camera pose and registers 3D features with metrics-aware covariance. In the experiments, MAC-VO outperforms existing VO algorithms even some SLAM algorithms in challenging environments.
Figure 2: MAC-VO pipeline. We use a shared matching network to estimate the depth, flow, and corresponding uncertainty. The learned uncertainty is leveraged to filter out unreliable features and model the metrics-aware covariance of the 3D keypoints. These registered keypoints and their covariance models are then utilized in the back-end optimization to determine the relative camera pose.
II Related Works
Existing geometric-based methods optimize the camera pose based on geometric constraints like re-projection error [1, 13] or photometric error [14, 15, 16]. To more accurately model the uncertainty of the depth, Civera et al. [17] and Montiel[18] investigate the inverse depth parameterization. These methods often use constant parameters and simple heuristics to model the covariance matrix of these errors during the factor graph optimization. For multi-sensor fusion [6, 19] and semantic SLAM [20, 21], the covariance model plays a significant role in weighing the confidence of different sensors and modules. To effectively capture sensor uncertainty, the covariance models are often tuned based on empirical prior. However, these simplified covariance models fail to capture the complexity of the challenging environments.
Recent advances in deep learning have transformed research in optical flow estimation [22, 23], feature matching [24, 10], depth estimation [25, 26, 27], and end-to-end camera pose estimation [28, 2]. Several methodologies have been developed to address the uncertainty in estimating depth, flow, and pose. Dexheimer et al. [29] proposed a learned depth covariance function that is applied in downstream tasks like 3D reconstruction. Nie et al. utilize a self-supervised learning method to jointly train depth and depth covariance of images in the wild [30]. ProbFlow [31] proposes using post-hoc confidence measures to assess the per-pixel reliability of flow.
Amidst these developments, hybrid learning-based SLAM systems [32, 26, 33, 34, 35, 36] are emerging to synergize the geometric constraint with the adaptability of learning-based methods. To improve the reliability of the feature tracking process, some methods introduce learning-based uncertainty measurements or confidence scores [37, 38, 10] for pose optimization. DROID-SLAM [3] utilizes a differentiable bundle adjustment layer to implicitly tune the uncertainty model. This method employs a simplified diagonal covariance model for the bundle adjustment. These methods focus on the relative confidence between features , but ignore the scale consistency of the covariance model.
For keypoint selection, geometric-based VO relies on hand-crafted features [8, 9], which detect the edges and corner points. The recent advance of the learning-based method train feature extractor in data-driven manner [7, 39]. However, these keypoint also prioritize the edge and corner features due to their data bias in the pre-training dataset. Recent works like D3VO [32] have shown that relying on edge and corner points can degrade state estimation, sometimes performing worse than random selectors [4]. The accuracy of learning-based feature matching and depth estimation algorithms is particularly compromised at object edges due to feature interpolation and the ambiguity in neural networks. In this work, we propose a keypoint selector based on learned uncertainty to filter out the unreliable features.
III Method
Figure 3: We include three filters: Non-minimum Suppression (NMS) filter, geometric filter, and uncertainty-based filter. In the KITTI dataset, the uncertainty filter implicitly filters out the unreliable feature matchings in the scene.
As illustrated in Figure 2, MAC-VO outlines an effective integration of a learning-based front-end and a geometrically constrained back-end using the metrics-aware covariance model. In the front-end (Section III-A), we train an uncertainty-aware matching network to model the corresponding uncertainty stems from feature deficiencies. Utilizing the learned uncertainty, we develop a keypoint selector in Section III-B to choose reliable features. The metrics-aware 2D uncertainty is then propagated to the 3D space via the proposed covariance model in Section. III-C. In the back-end optimization (Section. III-D), we optimize the relative motion by minimizing the distance between registered keypoints weighted by the 3D covariance.
Figure 4: Top: Architecture of the uncertainty-aware matching network. We employ a motion aggregator and an iterative update structure to enable the covariance module to capture inconsistencies in feature matching. Bottom: In each iteration, the model captures the inconsistency between the motions. For the $\Delta\sigma$, the red color indicates a positive $\Delta\sigma$ that increases the uncertainty, and blue means decreasing the uncertainty.
III-A Network & Uncertainty Training
The objective of our network is to predict the flow $\hat{f}\in\mathbb{R}^{2}$ and the corresponding uncertainty $\hat{\Sigma}_{f}=\mathrm{diag}(\sigma_{u}^{2},\sigma_{v}^{2})$. As shown in Fig. 4, an iterative motion aggregator inspired by the FlowFormer [11] is employed to capture the inconsistency in feature matching. To extend this network for uncertainty estimation, we add an uncertainty decoder to predict the $\Delta\sigma$ updates of the uncertainty in the log space. The use of log space enables additive updates and constrains the output range, stabilizing gradients and simplifying model output. After iterative updates, the log-variance passes through an exp activation function to obtain the final uncertainty. More details about the network are shown in Appendix. B-E.
To supervise the uncertainty, we leverage the negative log-likelihood loss used in conformal prediction [40, 41, 42]:
$\mathcal{L}_{cov}=\sum_{i}^{N}\alpha_{i}\left((y-f_{i})^{\top}\hat{\Sigma}_{f}^{-1}(y-f_{i})+\log(\det\hat{\Sigma}_{f})\right),$ (1)
where $y$ is the ground truth optical flow, $f_{i}$ denotes the $i$-th iteration of the network outputs, and $\alpha_{i}$ is the weight for each iteration and is set to decrease exponentially with ratio of $0.8$. During the training stage, we initialize the encoder network parameters with the pre-trained model by FlowFormer. We then train the covariance module on the synthetic dataset TartanAir [43]. Our experiments demonstrate that the model can generalize to real-world datasets without fine-tuning.
III-B Uncertainty-based Keypoint Selection
Different from the random selector used in DPVO[4] and the hand-crafted features used in ORB-SLAM [8], we leverage the learned uncertainty estimation to filter out unreliable features such as those on the vehicle illustrated in Fig. 3. This is achieved by composing three filters: the uncertainty filter, geometry filter, and the non-minimum suppression (NMS). The uncertainty filter removes pixels with depth and flow uncertainty larger than $1.5$ times the median uncertainty of the current frame, which discards the unreliable features while maintaining the diversity of keypoint candidates. The accurate uncertainty estimation effectively removes the keypoints on occluded objects, reflective surfaces, and feature-less areas. Illustrated in Fig. 3, the uncertainty filter removes all keypoint candidates on the moving vehicle on a KITTI trajectory due to its high flow uncertainty. Along with the uncertainty filter is the geometric filter, which constraints the disparity and removes keypoints on the edge of frame. To ensure the even distribution of keypoints in image, the NMS filter is applied on the product of depth and flow uncertainty map prior to both filters.
Figure 5: a) Uncertain estimation of depth due to the error in feature matching. b) Projecting depth and matching uncertainty from image plane to 3D space. c) Residual $\mathcal{L}_{i}$ for pose graph optimization.
III-C Metrics-aware 3D Covariance Model
In the context of the camera projection geometry, the covariance of a 3D keypoint is determined by the uncertainty of the depth $\sigma_{d}^{2}$ and matching $(\sigma_{u}^{2},\sigma_{v}^{2})$. To accurately model the covariance for 3D keypoint, it is critical to determine (1) the depth uncertainty of the matched points $\hat{\sigma}_{d}^{2}$ and (2) the off-diagonal covariance terms during the 2D-3D projection process.
Depth Uncertainty after keypoint matching  As shown in Fig. 5 (a), the matched features are expected to be within a probabilistic range centered at $[u_{i,t},v_{i,t}]$ due to flow uncertainty $(\sigma_{u_{i,t}}^{2},\sigma_{v_{i,t}}^{2})$. As a result, a minor disturbance in the feature matching may introduce a large error in the depth.
To address this problem, we model the depth uncertainty $\hat{\sigma}_{d_{i,t}}^{2}$ of the matched feature point based on the depth feature of the local patch $D_{i,t}$. We approximate it with the weighted sum of the variances within the patch. The weights are determined by a 2D Gaussian kernel $\varphi$, which utilizes $\sigma_{u_{i,t}}^{2}$ and $\sigma_{v_{i,t}}^{2}$ to adjust the influence of each point within the patch: $\sigma_{d_{i,t}}^{2}=\sum_{j}{\varphi_{j}(d_{j}-\mu_{{D}_{i,t}})^{2}}.$
Project 2D Uncertainty to 3D Covariance  Following the pinhole camera model with focus $\mathrm{f}_{x},\mathrm{f}_{y}$ and optical center $\mathrm{c}_{x},\mathrm{c}_{y}$, the coordinate of the keypoint is calculated by: $\mathbf{x}_{i,t}=(\mathbf{u}_{i,t}-\mathrm{c}_{x})\mathbf{d}_{i,t}/\mathrm{f}_{x}$, $\mathbf{y}_{i,t}=(\mathbf{v}_{i,t}-\mathrm{c}_{y})\mathbf{d}_{i,t}/\mathrm{f}_{y}$ and $\mathbf{z}_{i,t}=\mathbf{d}_{i,t}$, as shown in Fig. 5 (b). To accurately capture the uncertainties associated with these measurements, the main diagonal of the covariance matrix is formulated as:
$\sigma_{x_{i,t}}^{2}$ $={((\sigma_{u_{i,t}}^{2}+d^{2}_{i,t})(\sigma_{d_{i,t}}^{2}+u^{2}_{i,t})-u^{2}_{i,t}d^{2}_{i,t}+\mathrm{c}_{x}^{2}\sigma_{d_{i,t}}^{2})}/{\mathrm{f}_{x}^{2}},$ (2)
$\sigma_{y_{i,t}}^{2}$ $={((\sigma_{v_{i,t}}^{2}+d_{i,t}^{2})(\sigma_{d_{i,t}}^{2}+v_{i,t}^{2})-v_{i,t}^{2}d_{i,t}^{2}+\mathrm{c}_{y}^{2}\sigma_{d_{i,t}}^{2})}/{\mathrm{f}_{y}^{2}},$
$\quad\sigma_{z_{i,t}}^{2}$ $=\sigma_{d_{i,t}}^{2}.$
In this model, the projected coordinates are interdependent due to the common multiplier of depth $\mathbf{d}_{i,t}$. To precisely formulate the covariance of the 3D keypoints ${}^{c}\Sigma_{i,t}^{p}$ under camera coordinate, it is essential to include the off-diagonal covariance terms in ${}^{c}\Sigma_{i,t}^{p}$.
${}^{c}\Sigma_{i,t}^{p}=\begin{bmatrix}\sigma_{z}^{2}&\sigma_{xz_{i,t}}&\sigma_{yz_{i,t}}\\ \sigma_{xz_{i,t}}&\sigma_{x_{i,t}}^{2}&\sigma_{xy_{i,t}}\\ \sigma_{yz_{i,t}}&\sigma_{xy_{i,t}}&\sigma_{y_{i,t}}^{2}\\ \end{bmatrix},\;\begin{aligned} &\sigma_{xz}=\sigma_{d}^{2}(u-\mathrm{c}_{x})/\mathrm{f}_{x}\\ &\sigma_{yz}=\sigma_{d}^{2}(v-\mathrm{c}_{y})/\mathrm{f}_{y}\\ &\sigma_{xy}=\frac{\sigma_{d}^{2}(u-\mathrm{c}_{x})(v-\mathrm{c}_{y})}{\mathrm{f}_{x}\mathrm{f}_{y}}.\end{aligned}$ (3)
Results in the ablation study also confirm the necessity of off-diagonal terms for accurate pose graph optimization. Detailed derivation for Eq. 2 and Eq. 3 is in Appendix. A-C.
III-D Pose Graph Optimization
We optimize the camera pose $T_{t}\in SE(3)$ at time $t$ in the world frame by minimizing the distance of the matched 3D keypoints $p_{i,t-1}$ and ${}^{c}p_{i,t}$, where ${}^{c}p_{i,t}$ is in the camera frame. To reduce the initial error margin of the optimization, we initialize the camera pose using the relative motion estimated by the TartanVO [2].
The pose graph optimization is formulated as follows:
$T^{\star}$ $=\operatorname*{arg\,min}_{T_{t}}{\sum_{i}{\left\|p_{i,t-1}-{T}_{t}\,{}^{c}p_{i,t}\right\|^{2}_{\Sigma_{i}}}},$ (4)
$\Sigma_{i}$ $=\Sigma_{i,t-1}^{p}+R_{t}\;{}^{c}\Sigma_{i,t}^{p}\;R_{t}^{\top}.$
$\|\cdot\|_{\Sigma_{i}}$ represents the Mahalanobis distance with covariance matrix $\Sigma_{i}$. Unlike DROID-SLAM [3], which employs a diagonal covariance matrix, we model the correlation between axes to capture accurate inter-dependencies. We solved this optimization problem by Levenberg-Marquardt algorithms using PyPose [44].
IV Experiment
Datasets & Baseline  We evaluate the proposed model and baseline methods on a variety of public datasets, including synthetic dataset TartanAir v2 [43], real-world data from EuRoC [45], KITTI [46], as well as customized data collected from a Zed camera. These datasets cover a diverse range of hardware configurations, motion patterns, and environments. To demonstrate our method’s robustness under challenging scenarios, we collected the TartanAir v2, a new set of difficult trajectories following the TartanAir [43] that includes frequent indoor-outdoor transition and low-illumination scenes as shown in Fig. 6. To demonstrate the generalizability of our model, we use the same configuration across all datasets.
Evaluation Metrics  Our evaluation focuses more on relative error Since the proposed method does not contain loop closure or global bundle adjustment. So we use Relative translation error ($t_{\mathrm{rel}}$, $\mathrm{m/frame}$) and relative rotation error ($r_{\mathrm{rel}}$, $\mathrm{{}^{\circ}/frame}$) as:
$t_{\mathrm{rel}}$ $=\frac{1}{N}\sum^{N}_{t=1}{\left\|p_{t+1}-p_{t}-R_{t}\hat{R}_{t}^{\top}\!\!\left(\hat{p}_{t+1}-\hat{p_{t}}\right)\right\|_{2}},$ (5)
$r_{\mathrm{rel}}$ $=\frac{180}{\pi}\frac{1}{N}\sum^{N}_{t=1}\left\|\log\left(\hat{R}_{t,t+1}^{\top}R_{t,t+1}\right)\right\|_{2},$
where $p_{t}$ and $R_{t}$ are ground truth position and rotation, $\hat{p}_{t}$ and $\hat{R}_{t}$ is the estimated position and rotation. $R_{t,t+1}=R_{t}^{\top}R_{t+1}$ is the rotation from frame $t$ to frame $t+1$.
IV-A Quantitative Analysis
TABLE I: Performance comparison of different methods on the EuRoC Dataset. Only odd-ordered trajectory is shown due to page limit, see Appendix. B-A for the remaining results.
Trajectory MH01 MH03 MH05 V102 V201 V203 Avg.
$t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
SLAM
      ORB-SLAM 3 0.0035 0.0450 0.0058 0.0603 0.0059 0.0526 0.0096 0.1757 0.0064 0.1615 0.033 0.9497 0.0092 0.1866
      DROID-SLAM 0.0012 0.0159 0.0034 0.2656 0.0025 0.0193 0.0026 0.0417 0.0012 0.0289 0.0034 0.1033 0.0024 0.0590
VO
      TartanVO⋆† 0.0121 0.0560 0.0302 0.2791 0.0193 0.0604 0.0251 0.1244 0.0065 0.0920 0.0303 0.2986 0.0198 0.1270
      TartanVO 0.0277 0.5122 0.0514 0.6635 0.0464 0.4797 0.0394 1.0420 0.0195 0.4684 0.0473 1.9657 0.0368 0.8346
      iSLAM-VO 0.0042 0.0560 0.0076 0.2789 0.0070 0.0603 0.0066 0.1241 0.004 0.0920 0.0151 0.2984 0.0071 0.1269
      DPVO⋆† 0.0015 0.0207 0.0028 0.0273 0.0028 0.0243 0.0041 0.0496 0.0016 0.0342 0.0045 0.1205 0.0027 0.0437
Ours 0.0014 0.0214 0.0023 0.0238 0.0025 0.0216 0.0029 0.0434 0.0012 0.0289 0.0049 0.1284 0.0024 0.0403
Average over all trajectories of EuRoC.   Monocular method.   Scale-aligned with ground truth.
EuRoC Dataset    We assessed our model on the EuRoC [45] dataset, as detailed in Table. I, comparing it against baseline methods including visual odometries and state-of-the-art visual SLAM systems with loop-closure and global bundle adjustment. While our method exhibits compatible performance to DROID-SLAM on average $t_{\mathrm{rel}}$, it outperforms all baselines in terms of $r_{\mathrm{rel}}$ by around $10\%$.
TABLE II: Performance comparison on the TartanAir v2 Hard Dataset. Noticeably, ORB-SLAM3 lost track of all sequences and is therefore not included. Only the hard subset is shown due to page limit, see Appendix. VI for the remaining results.
Trajectory H00 H01 H02 H03 H04 H05 H06 Avg.
$t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
SLAM
      DROID-SLAM .0485 .1174 .0023 .0210 .0190 .0821 .0064 .0300 .0057 .0255 .1463 .2357 .0310 .0908 .0370 .0861
VO
      iSLAM-VO .4235 2.630 .3070 3.018 .3252 2.189 .3622 2.435 .2576 2.899 .2574 3.755 .2099 3.145 .3061 2.867
      TartanVO⋆† .1605 3.338 .2918 2.775 .2718 3.305 .2775 2.191 .2204 2.874 .1644 2.899 .2350 3.756 .2316 3.020
      TartanVO .1505 .4329 .0914 .7542 .0715 .3265 .0842 .4053 .0678 .6569 .0803 1.186 .0784 .9458 .0892 .6725
      DPVO⋆† .4984 .4937 .1738 .7112 .0539 .2539 .3847 2.703 .0481 .1869 .1891 1.430 .3365 2.943 .2406 1.246
Ours .0085 .1018 .0344 .1450 .0048 .0628 .0150 .0778 .0092 .1414 .0048 .0552 .0217 .4160 .0141 .1429
Monocular method.   Scale-aligned with ground truth.
TartanAir v2 Dataset    TartanAir v2 is challenging for visual SLAM. Our approach improves $61.9\%$ in $t_{\mathrm{rel}}$ compared to the nearest competitor. Notably, on trajectory H00, which simulates the lunar surface shown in Fig. 6, our model demonstrates a remarkable $82.4\%$ decrease in $t_{\mathrm{rel}}$ and achieves the lowest $r_{\mathrm{rel}}$ among all baseline methods.
TABLE III: Performance comparison of different methods on the KITTI Dataset. Only even numbered trajectory is shown due to page limit, see Appendix. VIII for remaining results.
Trajectory 00 02 04 06 08 10 Avg.
$t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
SLAM
      ORB-SLAM 3 0.0252 0.0586 0.0438 0.0529 0.0274 0.0322 0.0228 0.0338 0.0271 0.0455 0.0166 0.0515 0.0258 0.0434
      DROID-SLAM 0.0198 0.0538 0.0250 0.0445 0.0255 0.0296 0.0199 0.0296 0.0275 0.0381 0.0309 0.0715 0.0900 0.0448
VO
      TartanVO⋆† 0.2066 0.1055 0.1626 0.1105 0.1152 0.0789 0.2234 0.0816 0.1857 0.0823 0.1745 0.0907 0.2207 0.0886
      TartanVO 0.0656 0.1026 0.0905 0.1197 0.1747 0.1158 0.0923 0.0968 0.0721 0.1063 0.0679 0.0969 0.1804 0.1147
      iSLAM-VO 0.0577 0.1052 0.0686 0.1101 0.1356 0.0787 0.0837 0.0812 0.0510 0.082 0.0449 0.0905 0.0878 0.0883
      DPVO⋆† 0.4542 0.0495 0.4209 0.0381 0.0348 0.0219 0.2393 0.0250 0.3051 0.0347 0.0661 0.0386 0.1951 0.0329
Ours 0.0192 0.0654 0.0223 0.0715 0.0206 0.0473 0.0187 0.0456 0.0254 0.0509 0.019 0.0569 0.0420 0.0645
Average over all trajectories (from 00 to 10) of KITTI Odom.   Monocular method.   Scale-aligned with ground truth.
KITTI Dataset    To further validate our robustness and consistency in outdoor, large-scale trajectory with the presence of dynamic objects, we evaluate our system on the KITTI [46] dataset. Our method, which relies solely on two-frame pose optimization without incorporating multi-frame bundle adjustment or loop closure, shows commendable performance, ranked behind the ORB-SLAM3, a full visual SLAM system, on $t_{\mathrm{rel}}$. Our system significantly outperforms other visual odometry approaches in $t_{\mathrm{rel}}$, demonstrating a 53.3% reduction in relative translation error. The performance observed in $r_{\mathrm{rel}}$ may be attributed to the lack of multi-frame optimization.
IV-B Qualitative Analysis and Ablation Study
Figure 6: Top: Trajectories estimated by our model and baselines. We highlight the segments where dynamic objects interrupt the VO. Our method remains robust by implicitly filtering out unreliable features. Bottom left: We use our own platform with the ZED camera to collect data in the office. Our method demonstrates robustness against dynamic objects and visual occlusions in the images. Bottom right: samples of TartanAir v2 test dataset, which simulates the exotic lunar environment and factory with presence of dynamic objects.
In addition to quantitative evaluations, we conduct qualitative evaluations of our proposed system across multiple datasets including EuRoC, KITTI, and TartanAir v2, supplemented by manually collected data using a ZED stereo camera. Our model, even without multi-frame optimization, achieves top-tier performance and exhibits fewer glitches than baseline methods. As demonstrated in the Fig. 6, our method produces smoother trajectories and superior pose estimation precision. Fig. 6 a) presents that our model correctly identifies the region occluded by the mounted platform and dynamic objects in the scene and assigns a high uncertainty score to these regions.
Ablation Study   In Table. IV, we first perform the ablation study on each module of MAC-VO including (a. w/o CovKP & CovOPT) with random keypoint selector and identity covariance matrix. (b. w/o CovOpt) replace the metrics-aware covariance model with the identity covariance model. (c. w/o CovKP) replace the proposed keypoint selector with random keypoint selector.
To demonstrate the necessity of scale consistency and off-diagonal terms in our covariance model, we run the ablation study with different configurations: (I. DiagCov) remove off-diagonal terms in Eq. 3, (II. Scale-agnostic) normalize the covariance model by the average determinants of covariance matrices of each frame.
TABLE IV: Ablation study. Detailed results in Appendix. B-C.
Dataset TartanAir v2 Hard TartanAir v2 Easy
$t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
Module Ablation
      w/o CovKP & CovOpt .0743 .5221 .0521 .2799
      w/o CovOpt .0679 .3776 .0511 .2367
      w/o CovKP .0188 .2347 .0066 .0808
Covariance Ablation
      DiagCov .0461 .3023 .0277 .1548
      Scale Agnostic .0204 .2321 .0086 .0764
Ours .0141 .1429 .0051 .0670
Figure 7: Results of various ablation setups on the trajectory H01 of TartanAir v2.
IV-C Runtime Analysis
The runtime analysis shown in Table. V uses the platform with AMD Ryzen 9 5950X CPU and NVIDIA 3090 Ti GPU. We also introduce a fast mode (MAC-VO Fast) that utilizes half-precision number in the network inference to enhance efficiency. This mode also speeds up the memory decoder network by reducing the number of iterative updates from 12 to 4. The fast mode performs at 10.5 fps (frames per second) with 70% of the performance of the original MAC-VO. More details are included in Appendix. C.
TABLE V: Time spent on each module of MAC-VO under different optimization with image resolution of $640\times 640$.
Raw TRT MP + TRT MAC-VO Fast
Modules (ms)
     Frontend Network 401.9 239.3 239.9 81.6
     Optimization 53.4 57.5 - -
     Motion Model 5.2 5.1 5.3 5.2
     Keypoint Selector 0.7 0.6 0.6 0.5
     Covariance Model 0.6 0.7 0.7 0.7
Overall (fps) 2.15 3.25 3.96 10.57
TRT: TensorRT framework - https://developer.nvidia.com/tensorrt
MP: multi-processing the PGO in parallel with the matching network.
MAC-VO Fast: utilizes half-precision number (bf16) and light-weight model.
V Conclusion & Discussion
This paper proposes MAC-VO, a learning-based stereo VO method that leverages the metrics-aware covariance model. Our model outperforms most visual odometry and even SLAM algorithms on challenging datasets. In our current work, the model focuses on the two-frame pose optimization. We believe our accuracy will be further benefit from bundle adjustment, multi-frame optimization, and loop closure. Additionally, we are interested in exploring our metrics-aware covariance model in multi-sensor fusion, such as with IMUs.
Appendix A Covariance Model Formulation
A-A Depth Uncertainty From Disparity Uncertainty
In this section, we present the formulation for estimating the distribution of depth $\mathbf{d}$ on a single pixel given the estimated distribution of disparity ${\mathrm{Disp}}\sim\mathcal{N}(\mu_{{\mathrm{Disp}}},\sigma_{\mathrm{Disp}}^{2})$.
Following the pinhole camera model, the depth is calculated as $\mathbf{d}=bf_{x}/{\mathrm{Disp}}$ where the camera baseline is $b$. Since ${\mathrm{Disp}}$ can be zero, the distribution of $\mathbf{d}$ may be ill-defined. To fix this, we employ the first-order Taylor expansion to approximate $\mu_{d}$ and $\sigma_{d}^{2}$ such that $\mathbf{d}\sim\mathcal{N}(\mu_{d},\sigma_{d}^{2})$.
We assume the variance of disparity $\sigma_{{\mathrm{Disp}}}^{2}=(\gamma\mu_{\mathrm{Disp}})^{2}$ for some sufficiently small $\gamma>0$ such that the probability of ${\mathrm{Disp}}<=0$ is negligible. Based on this assumption, we have
$\mu_{\mathbf{d}}$ $=\mathbb{E}\Biggl{[}\frac{bf_{x}}{\mu_{\mathrm{Disp}}}+\left(\frac{d}{d\mu_{\mathrm{Disp}}}\frac{bf_{x}}{\mu_{\mathrm{Disp}}}\right)({\mathrm{Disp}}-\mu_{\mathrm{Disp}})\Biggr{.}$ (6)
+.$+\Biggl{.}\left(\frac{d^{2}}{d\mu_{\mathrm{Disp}}^{2}}\frac{bf_{x}}{\mu_{\mathrm{Disp}}}\right)\frac{({\mathrm{Disp}}-\mu_{\mathrm{Disp}})^{2}}{2}+\cdots\Biggr{]}$
$\approx\mathbb{E}\left[\frac{bf_{x}}{\mu_{\mathrm{Disp}}}\right]-\frac{1}{\mu_{\mathrm{Disp}}^{2}}\mathbb{E}[({\mathrm{Disp}}-\mu_{\mathrm{Disp}})]$
$=\frac{bf_{x}}{\mu_{\mathrm{Disp}}}-0=\frac{bf_{x}}{\mu_{\mathrm{Disp}}}$
Similarly, $\sigma_{\mathbf{d}}^{2}$ can be expressed as
$\sigma_{\mathbf{d}}^{2}$ $=\mathrm{Var}\Biggl{[}\frac{bf_{x}}{\mu_{\mathrm{Disp}}}+\left(\frac{d}{d\mu_{\mathrm{Disp}}}\frac{bf_{x}}{\mu_{\mathrm{Disp}}}\right)({\mathrm{Disp}}-\mu_{\mathrm{Disp}})\Biggr{.}$ (7)
+.$+\Biggl{.}\left(\frac{d^{2}}{d\mu_{\mathrm{Disp}}^{2}}\frac{bf_{x}}{\mu_{\mathrm{Disp}}}\right)\frac{({\mathrm{Disp}}-\mu_{\mathrm{Disp}})^{2}}{2}+\cdots\Biggr{]}$
$\approx\mathrm{Var}\left[\frac{bf_{x}}{\mu_{\mathrm{Disp}}}+\left(\frac{d}{d\mu_{\mathrm{Disp}}}\frac{bf_{x}}{\mu_{\mathrm{Disp}}}\right)({\mathrm{Disp}}-\mu_{\mathrm{Disp}})\right]$
$=\left(\frac{bf_{x}}{\mu_{\mathrm{Disp}}^{2}}\right)^{2}\cdot\sigma_{\mathrm{Disp}}^{2}=\frac{(bf_{x}\gamma)^{2}}{\mu_{\mathrm{Disp}}^{2}}$
Therefore, the approximation of depth uncertainty from disparity uncertainty is expressed as follows:
$D\sim\mathcal{N}\left(\frac{bf_{x}}{\mu_{\mathrm{Disp}}},\frac{(bf_{x}\gamma)^{2}}{\mu_{\mathrm{Disp}}^{2}}\right)$ (8)
Monte Carlo simulation indicates that for $\gamma<0.3$, the error of the aforementioned approximation is acceptable, as shown in Fig. 8.
Figure 8: Result of Monte Carlo simulation on depth distribution. Despite the skewness in the simulated distribution, our approximation matches the simulation. As error rate $\gamma$ increases and average disparity ${\mathrm{Disp}}$ approaches 00 (from right to left), the depth distribution becomes more skewed and the quality of approximation decreases.
A-B Depth Uncertainty with of Match Uncertainty
Let $q_{i,t}$ be the $i$-th keypoint on the camera plane at time $t$. Given the estimated optical flow $\mathbf{f}$ at $q_{i,t}$, the matched keypoint at time $t+1$ is defined as $q_{i,t+1}=q_{i,t}+\mathbf{f}_{i,t}$. Since $\mathbf{f}_{i,t}$ follows the gaussian distribution $\mathcal{N}(\hat{f}_{i,t},\hat{\Sigma}_{i,t})$, $q_{i,t+1}$ is a random variable following distribution of $\mathcal{N}(q_{i,t}+\hat{f}_{i,t},\hat{\Sigma}_{i,t})$.
Let $\varphi_{i,t}$ be a 2D Gaussian filter with covariance matrix $\hat{\Sigma}_{i,t}$, the probability for matched keypoint on some pixel $j$ is then expressed as $(\varphi_{i,t})_{j}$. Let $d_{j}$ denote the estimated depth at pixel $j$, then the average depth for pixel $q_{i,t+1}$ weighted by $\varphi$ is expressed as
$\mu_{d_{i,t}}=\sum_{j}{(\varphi_{i,t})_{j}\cdot d_{j}},$ (9)
and the estimated variance of depth of $q_{i,t+1}$ is calculated as weighted variance
$\sigma_{d_{i,t}}^{2}=\sum_{j}{(\varphi_{i,t})_{j}\cdot(d_{j}-\mu_{d_{i,t}})^{2}}.$ (10)
We could also model the depth of the matched point using a mixture of Gaussian distributions, but experiments show that this offers only a minimal performance improvement. Therefore, we use the straightforward weighted variance method to estimate depth uncertainty.
Figure 9: Estimate depth variance of the matched point under the presence of matching uncertainty
A-C Projecting 2D Uncertainty to Spatial Covariance
Let $\mathbf{u}_{i,t}\sim\mathcal{N}(u_{i,t},\sigma_{u_{i,t}}^{2})$, $\mathbf{v}_{i,t}\sim\mathcal{N}(v_{i,t},\sigma_{v_{i,t}}^{2})$, and $\mathbf{d}_{i,t}\sim\mathcal{N}(d_{i,t},\sigma_{d_{i,t}}^{2})$, we derive the distribution of 3D point under camera coordinate ${}^{c}\mathbf{p}_{i,t}=[\mathbf{x}_{i,t},\mathbf{y}_{i,t},\mathbf{z}_{i,t}]^{\top}\sim\mathcal{N}({}^{c}p_{i,t},{}^{c}\Sigma_{i,t})$.
Recall that the relationship between pixel coordinate $(\mathbf{u}_{i,t},\mathbf{v}_{i,t})$, depth $\mathbf{d}_{i,t}$ and 3D coordinate $\mathbf{x}_{i,t},\mathbf{y}_{i,t},\mathbf{z}_{i,t}$ is depicted as
$\mathbf{x}_{i,t}=\frac{(\mathbf{u}_{i,t}-c_{x})\mathbf{d}_{i,t}}{f_{x}},\quad\mathbf{y}_{i,t}=\frac{(\mathbf{v}_{i,t}-c_{y})\mathbf{d}_{i,t}}{f_{y}},\quad\mathbf{z}_{i,t}=\mathbf{d}_{i,t}$ (11)
Assume $\mathbf{u}_{i,t}$, $\mathbf{v}_{i,t}$, $\mathbf{d}_{i,t}$ are independent to each other, we have $\mathrm{Var}(\mathbf{u}_{i,t},\mathbf{d}_{i,t})=(\sigma_{u_{i,t}}^{2}+d_{i,t}^{2})(\sigma_{d_{i,t}}^{2}+u_{i,t}^{2})-u_{i,t}^{2}d_{i,t}^{2}$ [47]. Based on this expression of variance, it follows that
$\sigma_{x_{i,t}}^{2}$ $=\mathrm{Var}\left(\frac{\mathbf{u}_{i,t}\mathbf{d}_{i,t}}{f_{x}}-\frac{c_{x}\mathbf{d}_{i,t}}{f_{x}}\right)=\frac{\mathrm{Var}(\mathbf{u}_{i,t}\mathbf{d}_{i,t})}{f_{x}^{2}}+\frac{c_{x}^{2}\mathrm{Var}(\mathbf{d}_{i,t})}{f_{x}^{2}}$ (12)
$=\frac{(\sigma_{u_{i,t}}^{2}+d_{i,t}^{2})(\sigma_{d_{i,t}}^{2}+u_{i,t}^{2})-u_{i,t}^{2}d_{i,t}^{2}+c_{x}^{2}\sigma_{d_{i,t}}^{2}}{f_{x}^{2}}$
$\sigma_{y_{i,t}}^{2}$ $=\mathrm{Var}\left(\frac{\mathbf{v}_{i,t}\mathbf{d}_{i,t}}{f_{y}}-\frac{c_{y}\mathbf{d}_{i,t}}{f_{y}}\right)=\frac{\mathrm{Var}(\mathbf{v}_{i,t}\mathbf{d}_{i,t})}{f_{y}^{2}}+\frac{c_{y}^{2}\mathrm{Var}(\mathbf{d}_{i,t})}{f_{y}^{2}}$
$=\frac{(\sigma_{v_{i,t}}^{2}+d_{i,t}^{2})(\sigma_{d_{i,t}}^{2}+v_{i,t}^{2})-v_{i,t}^{2}d_{i,t}^{2}+c_{y}^{2}\sigma_{d_{i,t}}^{2}}{f_{y}^{2}}$
$\sigma_{z_{i,t}}^{2}$ $=\sigma_{d_{i,t}}^{2}$
Under the assumption that $\mathbf{u}_{i,t},\mathbf{v}_{i,t},\mathbf{d}_{i,t}$ are independent to each other, we derive the covariance between $\mathbf{x}_{i,t}$, $\mathbf{y}_{i,t}$ and $\mathbf{z}_{i,t}$ as:
$\mathrm{Cov}(\mathbf{x}_{i,t},\mathbf{y}_{i,t})$ $=\mathrm{Cov}\left(\frac{(\mathbf{u}_{i,t}-c_{x})\mathbf{d}_{i,t}}{f_{x}},\frac{(\mathbf{v}_{i,t}-c_{y})\mathbf{d}_{i,t}}{f_{y}}\right)$ (13)
$=\mathbb{E}\left[\frac{\mathbf{d}_{i,t}^{2}(\mathbf{u}_{i,t}-c_{x})(\mathbf{v}_{i,t}-c_{y})}{f_{x}f_{y}}\right]$
$-\mathbb{E}\left[\frac{(\mathbf{u}_{i,t}-c_{x})\mathbf{d}_{i,t}}{f_{x}}\right]\mathbb{E}\left[\frac{(\mathbf{v}_{i,t}-c_{y})\mathbf{d}_{i,t}}{f_{y}}\right]$
$=\frac{(\mathbb{E}[\mathbf{d}_{i,t}^{2}]-\mathbb{E}[\mathbf{d}_{i,t}]^{2})\mathbb{E}[\mathbf{u}_{i,t}-c_{x}]\mathbb{E}[\mathbf{v}_{i,t}-c_{y}]}{f_{x}f_{y}}$
$=\frac{\sigma_{d_{i,t}}^{2}}{f_{x}f_{y}}(u_{i,t}-c_{x})(v_{i,t}-c_{y})$
and
$\mathrm{Cov}(\mathbf{x}_{i,t},\mathbf{z}_{i,t})$ $=\mathrm{Cov}\left(\frac{(\mathbf{u}_{i,t}-c_{x})\mathbf{d}_{i,t}}{f_{x}},\mathbf{d}_{i,t}\right)$ (14)
$=\frac{(\mathbb{E}[\mathbf{d}_{i,t}^{2}]-\mathbb{E}[\mathbf{d}_{i,t}]^{2})\mathbb{E}[\mathbf{u}_{i,t}]}{f_{x}}=\frac{\sigma_{d_{i,t}}^{2}}{f_{x}}(u_{i,t}-c_{x})$
$\mathrm{Cov}(\mathbf{y}_{i,t},\mathbf{z}_{i,t})$ $=\mathrm{Cov}\left(\frac{(\mathbf{v}_{i,t}-c_{y})\mathbf{d}_{i,t}}{f_{y}},\mathbf{d}_{i,t}\right)$
$=\frac{(\mathbb{E}[\mathbf{d}_{i,t}^{2}]-\mathbb{E}[\mathbf{d}_{i,t}]^{2})\mathbb{E}[\mathbf{v}_{i,t}]}{f_{y}}=\frac{\sigma_{d_{i,t}}^{2}}{f_{y}}(v_{i,t}-c_{y})$
Fig. 10 visualize the distribution of keypoints in 3D space via Monte Carlo and the 90% confidence interval of estimated distribution, confirming the necessity of off-diagonal terms.
Figure 10: Comparison of proposed 3D covariance (left) and the diagonal covariance matrix (right, DiagCov in ablation study). Our method captures the uncertainty of keypoints significantly better than the diagonal covariance matrix.
Appendix B Additional Results
B-A Remaining Results on EuRoC, TartanAirv2, and KITTI
TABLE VI: Performance comparison of different methods on the EuRoC Dataset. Only even-ordered trajectory is shown here, see Table. I for the remaining results.
Trajectory MH02 MH04 V101 V103 V202
$t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
SLAM
      ORB-SLAM 3 0.0036 0.0495 0.0061 0.0501 0.0049 0.0888 0.0137 0.2669 0.0090 0.1528
      DROID-SLAM 0.0012 0.0169 0.0031 0.0224 0.0024 0.0314 0.0036 0.0642 0.0017 0.0399
VO
      TartanVO⋆† 0.0172 0.0621 0.0213 0.0681 0.0124 0.0756 0.0263 0.1552 0.0171 0.1251
      TartanVO 0.0289 0.5037 0.0501 0.5400 0.0224 0.5322 0.0351 1.3127 0.0361 1.1607
      iSLAM-VO 0.0041 0.0620 0.0082 0.0682 0.0041 0.0756 0.0088 0.1554 0.0078 0.1252
      DPVO⋆† 0.0014 0.0212 0.0029 0.0264 0.0026 0.0405 0.0033 0.0662 0.0022 0.0493
Ours 0.0013 0.0199 0.0028 0.0273 0.0024 0.0304 0.0032 0.058 0.0018 0.0406
The estimated sequence is scale-aligned with ground truth.
Monocular method.
TABLE VII: Relative motion error on the TartanAir v2 Hard Dataset. Only even numbered trajectories in the Hard subset are shown here. For odd-numbered trajectories in the Hard subset, see Table. VIII. ORB-SLAM3 lost track of all trajectories and is therefore not shown in the table.
Trajectory H00 H02 H04 H06 H08 H10 H12
$t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
SLAM
      DROID-SLAM 0.009 0.033 0.029 0.080 0.014 0.026 0.006 0.025 0.019 0.082 0.004 0.060 0.146 0.236
VO
      iSLAM-VO 0.182 3.337 0.341 2.557 0.422 2.776 0.382 2.423 0.325 2.189 0.259 2.874 0.257 3.755
      TartanVO⋆† 0.097 0.459 0.315 2.629 0.247 3.084 0.234 3.017 0.272 3.305 0.245 2.435 0.164 2.899
      TartanVO 0.067 0.751 0.070 0.382 0.150 0.724 0.106 0.398 0.072 0.327 0.082 0.809 0.080 1.186
      DPVO⋆† 0.010 0.077 0.025 0.079 0.308 0.654 0.024 0.066 0.054 0.254 0.164 0.250 0.189 1.430
Ours 0.011 0.120 0.008 0.058 0.033 0.168 0.017 0.090 0.016 0.085 0.011 0.162 0.027 0.604
The estimated sequence is scale-aligned with ground truth.
Monocular method.
TABLE VIII: Performance comparison on the TartanAir v2 Easy Dataset.
Trajectory E00 E01 E02 E03 E04 E05 E06 Avg.
$t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
SLAM
      ORB-SLAM3 .1019 2.349
      DROID-SLAM .0077 .0144 .0025 .0199 .0063 .0409 .0049 .0251 .0009 .0147 .0031 .0463 .0016 .0235 .0039 .0264
VO
      iSLAM-VO .0656 .2873 .0456 .3853 .0359 .2234 .0508 .2635 .0268 .3201 .0464 .6624 .0362 .4606 .0439 .3718
      TartanVO⋆† .0532 .3237 .0937 .4750 .1066 .6048 .0756 .2230 .1114 .4032 .0862 .3185 .1373 .6620 .0949 .4300
      TartanVO .0505 .1334 .0322 .2078 .0237 .1105 .0303 .1417 .0173 .1681 .0279 .3828 .0218 .2329 .0291 .1967
      DPVO⋆† .0113 .0187 .0047 .0249 .0099 .0475 .0603 .2064 .0044 .0177 .0511 .0665 .0189 .1543 .0229 .0766
Ours .0026 .0351 .0124 .1183 .0031 .0684 .0054 .0383 .0050 .0413 .0018 .0247 .0054 .1427 .0051 .0670
Monocular method.   Scale-aligned with ground truth.
TABLE IX: Performance comparison of different methods on the KITTI Dataset. Only odd-numbered trajectory is shown here, see Table. III for the remaining results.
Trajectory 01 03 05 07 09
$t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
SLAM
      ORB-SLAM 3 0.0416 0.0355 0.027 0.0425 0.0161 0.0416 0.0155 0.0385 0.0208 0.0444
      DROID-SLAM 0.7112 0.0406 0.0182 0.0385 0.0153 0.0353 0.0746 0.0734 0.0214 0.0378
VO
      TartanVO⋆† 0.6834 0.0895 0.1234 0.0682 0.1821 0.0761 0.2005 0.0847 0.1704 0.1069
      TartanVO 1.1408 0.2455 0.0477 0.0953 0.0637 0.0821 0.0700 0.0931 0.0990 0.1077
      iSLAM-VO 0.2978 0.0896 0.0507 0.0681 0.0504 0.0758 0.0593 0.0842 0.0660 0.1064
      DPVO⋆† 0.0942 0.0247 0.0302 0.0330 0.2221 0.0319 0.1064 0.0311 0.1723 0.0336
Ours 0.1670 0.1670 0.0504 0.0504 0.0466 0.0466 0.0507 0.0507 0.0567 0.0567
Average is calculated over all trajectories (from 00 to 10) of KITTI.
The estimated sequence is scale-aligned with ground truth.
Monocular method.
B-B Robustness Analysis
TABLE X: Robustness of systems on TartanAir v2 Hard test dataset demonstrate by the average variance of relative translation and rotation error.
Model DROID-SLAM iSLAM-VO TartanVO⋆† TartanVO DPVO⋆† Ours
Avg. $\sigma^{2}_{t_{\mathrm{rel}}}$ 0.072 0.383 0.169 0.107 0.318 0.045
Avg. $\sigma^{2}_{r_{\mathrm{rel}}}$ 0.418 2.925 2.748 0.991 2.099 0.475
The estimated sequence is scale-aligned with ground truth.
Monocular method.
B-C Additional Ablation Study
TABLE XI: Performance comparison of different ablation setups on the TartanAir v2 Hard Dataset.
Relative Translation Error ($t_{\mathrm{rel}},\mathrm{m}/\mathrm{frame}$)
Trajectory H00 H01 H02 H03 H04 H05 H06
System Modules
w/o CovKP & CovOpt 0.136 0.037 0.058 0.098 0.027 0.071 0.092
w/o CovOpt 0.171 0.033 0.061 0.106 0.023 0.034 0.048
w/o CovKP 0.025 0.005 0.015 0.025 0.005 0.029 0.028
Covariance Model
DiagCov 0.128 0.016 0.038 0.066 0.014 0.023 0.038
Scale Agnostic 0.041 0.004 0.018 0.027 0.005 0.021 0.027
Ours 0.008 0.034 0.005 0.015 0.009 0.005 0.020
Relative Rotation Error ($r_{\mathrm{rel}},{}^{\circ}/\mathrm{frame}$)
Trajectory H00 H01 H02 H03 H04 H05 H06
System Modules
w/o CovKP & CovOpt 0.380 0.277 0.206 0.456 0.248 0.926 1.160
w/o CovOpt 0.350 0.244 0.248 0.507 0.176 0.449 0.669
w/o CovKP 0.099 0.042 0.076 0.247 0.059 0.629 0.491
Covariance Model
DiagCov 0.333 0.126 0.192 0.404 0.137 0.379 0.544
Scale Agnostic 0.173 0.042 0.098 0.253 0.060 0.500 0.498
Ours 0.102 0.145 0.063 0.078 0.141 0.055 0.465
TABLE XII: Performance comparison of different ablation setups on the TartanAir v2 Easy Dataset.
Relative Translation Error ($t_{\mathrm{rel}},\mathrm{m}/\mathrm{frame}$)
Trajectory E00 E01 E02 E03 E04 E05 E06
System Modules
w/o CovKP & CovOpt 0.116 0.049 0.045 0.058 0.018 0.035 0.044
w/o CovOpt 0.124 0.052 0.048 0.070 0.016 0.020 0.027
w/o CovKP 0.010 0.003 0.006 0.008 0.002 0.009 0.009
Covariance Model
DiagCov 0.069 0.025 0.025 0.037 0.009 0.012 0.017
Scale Agnostic 0.026 0.004 0.006 0.009 0.002 0.005 0.008
Ours 0.003 0.012 0.003 0.005 0.005 0.002 0.005
Relative Rotation Error ($r_{\mathrm{rel}},{}^{\circ}/\mathrm{frame}$)
Trajectory E00 E01 E02 E03 E04 E05 E06
System Modules
w/o CovKP & CovOpt 0.238 0.249 0.151 0.204 0.152 0.492 0.473
w/o CovOpt 0.219 0.254 0.187 0.275 0.132 0.278 0.313
w/o CovKP 0.048 0.031 0.039 0.058 0.025 0.189 0.174
Covariance Model
DiagCov 0.189 0.127 0.106 0.167 0.085 0.179 0.210
Scale Agnostic 0.073 0.033 0.027 0.067 0.029 0.148 0.136
Ours 0.035 0.118 0.068 0.038 0.041 0.025 0.143
B-D Datasets and Implementation
TartanAir dataset
The Tartanair dataset is a large-scale synthetic dataset encompassing highly diverse scenes, including various complex and challenging environments. Following the TartanAir data generation method, we created new, more diverse and challenging trajectories. From these, we selected that feature fast camera movements, low-light indoor environments, and simulated lunar surfaces lacking visual features. The images in the new dataset have a resolution of $640\times 640$ When testing iSLAM-VO and TartanVO, we resized the input images to $448\times 640$ to match the input requirements of the optical flow network. Due to the substantial GPU memory required by DROID-SLAM for global bundle adjustment optimization when processing high-resolution images, we reduced the input image resolution to $512\times 512$ and manually reclaimed GPU and memory after testing each trajectory to avoid potential memory insufficiency.
KITTI dataset
The KITTI dataset is a well-known and widely used dataset for autonomous driving, containing detailed ground truth labels that make it suitable for evaluating the performance of various VO/SLAM methods. Learning-based methods may experience performance degradation when handling the image sizes in the KITTI dataset, therefore, we cropped the input images to different sizes based on the methods tested. Our model processed images cropped to $376\times 680$ , while for testing DROID-SLAM, the images were cropped to $320\times 832$
EuRoC dataset
The Euroc dataset consists of 11 trajectories collected by a drone and is also a widely used benchmark for VO/SLAM tasks. Some scenes in this dataset contain thousands of image pairs, which imposes computational pressure on DROID-SLAM when performing bundle adjustment. Thus, during testing, we reduced the input image resolution to $320\times 512$ and read every other frame of image pairs. In post-processing, we completed the entire trajectory by interpolating timestamps.
ZED Camera data
The ZED Camera data comprises real-world trajectory data captured using a ZED stereo camera, employed to test the robustness and generality of our model when faced with unseen data not present in the training set.
B-E Network and Training Details
In the uncertainty-aware matching network, we employ a flow decoder similar to Flowformer to obtain $\Delta f$. In each iteration, the flow $f_{x}$ is updated by $f_{x}\leftarrow f_{x}+\delta f$. With the shared features, we train a new decoder with ConvGRU layers to estimate the uncertainty updates $\Delta\Sigma$. To stabilize the training, we employ $\exp$ as the activation function of the covariance decoder.
For each pixel $x$, we extract the local cost-map patch $q_{x}$ from the cost map, and the context feature $t_{x}$ from the context encoder. We encode the $q_{x}$ using a transformer FFN. Given the 2D position $p=x+f_{x}$, we encode it into positional embedding $\text{PE}(p)$. We aggregate these features using CNN encoder ME to generate local motion features ${I_{x}}$. Utilizing the GMA module, the network’s global motion features $G_{x}$ are acquired from the current motion features $\text{Att}_{f}(t_{x})$ and ${I_{x}}$. Subsequently, the shared motion features $m_{x}$ is obtained by concatenating the $G_{x}$, $I_{x}$ and $\text{Att}(t_{x})$, where $\text{Att}(t_{x})$ is the attention of the context features. To estimate the flow update $\Delta f$ and $\Delta\Sigma$, we us the ConvGRU $\Sigma_{\pi}(m_{x})$
$I_{x}$ $=\text{ME}(f_{x},\text{Concat}(\text{FFN}(q_{x}),\text{PE}(p)))$ (15)
$G_{x}$ $=\text{GMA}(\text{Att}(t_{x}),{I_{x}})$
$m_{x}$ $=\text{Concat}(\text{Att}(t_{x}),{I_{x}},G_{x})$
$\Delta f$ $=\text{FlowGRU}(m_{x})$
$\Delta\Sigma$ $=\text{CovGRU}(m_{x})$
We use AdamW optimizer with a learning rate of $12.5\times 10^{-5}$. The model is trained on A100 GPU consuming 16 GB GPU memory.
B-F GPU memory & Parameters
As shown in Table. XIII, we use 4.20GB of GPU memory, which is 6.7 times smaller than DROID-SLAM. This is because we utilize the sparse features for back-end optimization, reducing the requirements for the GPU memory.
TABLE XIII: The GPU memory consumption and the number of parameters during testing
Methods GPU memory Parameters
Ours 4.20GB 19.2M
TartanVO-stereo 1.16GB 47.3M
DROID-SLAM 28.34GB 4.0M
DPVO 1.51GB 3.4M
Appendix C Runtime Analysis
Figure 11: Multiprocessing setup places the optimizer in the background process and parallelizes the CPU-intensive (optimizer) and GPU-intensive (frontend network inference) jobs.
The runtime analysis use the platform with AMD Ryzen 9 5950X CPU and NVIDIA 3090 Ti GPU. To speed up the frontend network, we utilize CUDAGraph [^1] https://docs.nvidia.com/cuda/cuda-c-programming-guide/index.html#cuda-graphs (CUDAGraph), TensorRT [^2] https://developer.nvidia.com/tensorrt (TRT), and multi-processing (MP) to accelerate the original network (Raw). In our multiprocessing setup, as shown in Fig. B, a new CPU process is initiated to run the pose graph optimization in parallel with the front-end network, thereby maximizing GPU utilization.
We also introduce a fast mode (MAC-VO Fast) which utilizes half-precision (float16) number in the network inference to enhance computational efficiency. This mode also speeds up the memory decoder network by reducing the number of iterative updates from 12 to 4. The fast mode performs 10.5 fps (frames per second) with 70% of the performance of the original MAC-VO.